OpenStudy (marigirl):

Related Rates of change- Differentiation Show that:

OpenStudy (marigirl):

Show that: \[\frac{ d }{ dx }(\frac{ v^2 }{ 2 })=\frac{ d^2x }{ dt^2 } \] Where x represents the distance traveled in t secs by a particle moving with velocity v

OpenStudy (marigirl):

\[\frac{ dx }{ dt }=v\] Could we state that?

OpenStudy (baru):

hi

OpenStudy (baru):

yes dx/dt is v

OpenStudy (baru):

differentiate dx/dt=v again with respect to time to get \[d^2x/dt^2=dv/dt\]

OpenStudy (baru):

now you can replace the right hand side of the given equation with "dv/dt". now all thats left to do, is to show that the left hand side of the equation some how reduces to dv/dt. with me so far?

OpenStudy (baru):

do you know the "chain rule" of differentiation?

OpenStudy (marigirl):

oh yup got it! thanks!